range of the function f(x) = 5x − 3 defined by the domain {0,1,2}


1
Expert's answer
2021-12-14T04:51:57-0500
"f(0)=5(0)-3=-3"

"f(1)=5(1)-3=2"

"f(2)=5(2)-3=7"

Range: "\{-3, 2, 7\}"